Somebody solve this please........the length of a rectangular mailing label is 5 centimeters less than twice the width. The perimeter is 38 centimeters. Find the dimensions of the label......
Answer by jim_thompson5910(35256)




Let x be the width


"The length of a rectangular mailing label is 5 centimeters less than twice the width" means that the length is 2x-5


We have these three facts:
W = x
L = 2x-5
P = 38


We'll use those three facts above to find x. The formula used is the rectangle perimeter formula


P = 2*(L+W)


38 = 2*(x+2x-5) ... substitute (see equations above)


38 = 2*(3x-5)


38 = 2*(3x)+2(-5)


38 = 6x-10


6x-10 = 38


6x-10+10 = 38+10


6x = 48


6x/6 = 48/6


x = 8


--------------------


We know now that x = 8. So the width is 8 cm. Recall that at the top we let x = width.


The length is L = 2x-5 = 2*8-5 = 16-5 = 11 cm


--------------------


In summary, the dimensions are 11 cm by 8 cm